OpenStudy (anonymous):

An unknown compound contains only C, H, and O. Combustion of 3.10 g of this compound produced 7.05 g of CO2 and 2.88 g of H2O.

OpenStudy (aaronq):

apply the exact same concept as the last question.

OpenStudy (aaronq):

except this time you also have oxygen

OpenStudy (anonymous):

i don't fully understand what you did last time...

OpenStudy (aaronq):

find the moles of each moles of carbon = moles of CO2 moles of H are = moles of H2O (x2) (because there are 2 H atoms) moles of O are = moles of H2O + moles of CO2 (x2) .. again, because 2 atoms of O.

OpenStudy (aaronq):

divide them all by the one with the least moles, and multiply them by a number to get them all as whole numbers.

OpenStudy (anonymous):

so 7.05/44=0.160227273/0.159857904=1.002310607 and 2.88/18.016=0.159857904/0.159857904=1 this makes no sense..

OpenStudy (toxicsugar22):

hey after you helap her canone or both of you guys help me on my question i posted in chem

OpenStudy (aaronq):

do it in steps; moles of carbon= 7.05/44=0.160227273 moles of H 2.88/18.016 *2= moles of O = 0.160227273*2+(2.88/18.016)=

OpenStudy (anonymous):

2.88/18.016 *2=0.319715808 0.160227273*2+(2.88/18.016)=0.48031245 now what?

OpenStudy (aaronq):

divide all by the species with the lowest amount of moles

OpenStudy (anonymous):

so 0.160227273/0.160227273=1 0.319715808/0.160227273=1.995389437=2 0.48031245/0.160227273=2.997694718=3 so its \[CH _{2} O _{3}\]?

OpenStudy (anonymous):

nope.. that's not it.. ): i actually got excited for a minute.. I suck..

OpenStudy (anonymous):

\[CH _{9}O _{6}\]?

OpenStudy (anonymous):

dang it! nope.. that's not it either..

OpenStudy (anonymous):

help me.. i've tried four times and was wrong everytime..

OpenStudy (anonymous):

that's it.. i give up..

OpenStudy (anonymous):

\[C _{3}H _{6}O\] is the answer..

OpenStudy (anonymous):

thanks for all your help though!